# Failure of the Mott’s formula for the Thermopower in Carbon Nanotubes

### v1

@@ Abstract
$\sigma^{(sm)}\!=\!\frac{e^{2}}{4h}\frac{1}{4T}\int_{-\infty}^{\infty}\left[3\!+\!\frac{2}{\pi}\arctan\left(\left(E\!-\!E_{p}\right)\tau\right)\right]\frac{dE}{\cosh^{2}\frac{E\!-\!\mu}{2T}}.$ (23)
Integration of the second term by parts followed by the application of the Cauchy’s theorem (see Supplemental Material, Sec. S3 [14]) yields the conductivity explicitly:
$\sigma^{(sm)}\!=\!\frac{{{e^{2}}}}{{4h}}\left[{3\!-\!\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}+\frac{1}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}\right].$ (24)
@@
$S_{M}^{\left({{{sm}}}\right)}(\mu)=\frac{1}{{3e}}\,\frac{{{\mathop{\rm Re}\nolimits}{\psi^{\prime}}\left({\frac{1}{2}+\frac{1}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}}{{3-\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}+\frac{1}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}}.$ (25)
@@
In the limiting cases when the chemical potential is either close to the bottom of the parabolic subband or sufficiently far from it one finds
$S_{M}^{(sm)}=\frac{1}{{3e}}\begin{cases}\frac{1}{3}{\mkern 1.0mu}{\psi^{\prime}}\left({\frac{1}{2}+\frac{1}{{2\pi{\mkern 1.0mu}T\tau}}}\right)+\left[\psi^{\prime}\left({\frac{1}{2}+\frac{1}{{2\pi{\mkern 1.0mu}T\tau}}}\right)\right]^{2}\cdot\frac{{{E_{p}}-\mu}}{72\pi^{2}T}\\ \qquad\qquad\qquad\qquad\text{when}\qquad|E_{p}-\mu|\ll T,\\ \frac{{32\pi}}{11\cdot T\tau}\frac{{T^{2}}}{{{{\left({{E_{p}}-\mu}\right)}^{2}}}}\qquad\text{when}\qquad|E_{p}-\mu|\gg T.\end{cases}.$
@@
Figure 4 (a,b) shows the results of calculation of the Seebeck coefficient as a function of the chemical potential in the vicinity of the bottom of the second subband performed accounting for the smearing of DOS. The curves in the panels (a) and (b) are calculated with use of the Kelvin’s and Mott’s approaches, respectively. One can see that even in the case of a strong smearing Kelvin’s and Mott’s approaches yield qualitatively different results, especially in what concerns the change of sign of the Seebeck coefficient in the vicinity of the topological phase transition point.
@@
In conclusion, the ballistic regime offers peculiar modifications of some well-known thermoelectric relations. In particular, it turns out that the Mott’s formula for the Seebeck effect is incorrect for one-dimensinal parabolic bands, while the Kelvin’s formula remains fully accurate in this case. The differences between two approaches become apparent in the vicinity of resonances between the electronic chemical potential and the bottoms of energy subbands characterised by parabolic dispersions. The Mott’s formula predicts a finite value of the Seebeck coefficient at the resonance and no change of sign. In contrast, the Kelvin’s formula predicts divergence and the change of sign of the Seebeck coefficient at the resonance between the chemical potential and the bottom of a parabolic subband. The latter result is characteristic of most of topological phase transitions that occur once a new energy subband comes into play. Our analysis shows that both Kelvin’s and Mott’s expressions predict zero contributions to the Seebeck coefficient from the linear dispersion band in metallic CNTs in the ballistic regime. In contrast, the contribution of parabolic bands is not zero even far from the resonance points, according to the Kelvin’s (but not Mott’s!) formula. This conclusion can be easily verified experimentally. Any deviation of the Seebeck coefficient of CNTs from zero in the ballistic regime would characterize the inaccuracy of the Mott’s formula. Finally, we note that in 2D and 3D cases, the Kelvin’s formula also fails in the ballistic regime. This is because counter-propagating non-dissipative currents of ”hot” and ”cold” electrons are formed due to the combined actions of the temperature and voltage drops in these cases. The total electric current remains zero, but the system cannot be described by a single electro-chemical potential. We are looking forward for the experimental manifestations of these theoretical results.
Supplemental material for “Failure of the Mott’s formula for the Thermopower in Carbon Nanotubes”
@@
I S1: The link between the temperature derivative of the chemical potential and the density of states
Indeed, the temperature derivative of the chemical potential can be expressed as follows:
$\left(\frac{\partial\mu}{\partial T}\right)_{n}=\left(\frac{\partial n}{\partial T}\right)_{\mu}\left(\frac{\partial n}{\partial\mu}\right)_{T}^{-1}.$ (S1)
@@ I S1: The link between the temperature derivative of the chemical potential and the density of states
$n\left(\mu,T\right)=\int\limits_{0}^{+\infty}\frac{\nu(E)dE}{\exp\left(\frac{E-\mu}{T}\right)+1}.$ (S2)
Required derivatives can be easily obtained in the general form
$\left(\frac{\partial n}{\partial T}\right)_{\mu}=\frac{\pi^{2}}{3}T\frac{d\nu\left(\mu\right)}{d\mu}.$ (S3)
@@
$\left(\frac{\partial n}{\partial\mu}\right)_{T}=\nu\left(\mu\right).$ (S4)
Hence
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)=\frac{\pi^{2}}{3e}T\frac{\partial}{\partial\mu}\ln\nu(\mu).$ (S5)
II S2: The vicinity of the Dirac point
@@
$n_{e,h}\left(\mu,T\right)=\frac{2}{\pi v_{F}}\int\limits_{0}^{+\infty}\frac{dE}{\exp\left(\frac{E-\mu_{e,h}}{T}\right)+1},$ (S6)
@@
$n_{e}-n_{h}=n_{0}.$ (S7)
Here $-|e|n_{0}$ is the overall charge density. Substituting Eq. (S6) in Eq. (S7) and performing integration one finds
$\mu=\mu_{e}=\frac{\pi v_{F}n_{0}}{2}.$ (S8)
@@
$S_{K}=\frac{1}{e}\left(\frac{d\mu}{dT}\right)=\frac{1}{e}\left(\frac{\partial n_{0}}{\partial T}\right)_{\mu}\left(\frac{\partial n_{0}}{\partial\mu}\right)_{T}^{-1}=0.$ (S9)
II.1 The special point $\mu=0$
@@
$\left[\left(\frac{\partial n_{e}}{\partial T}\right)_{\mu_{e}}-\left(\frac{\partial n_{h}}{\partial T}\right)_{\mu_{h}}\right]\left[\left(\frac{\partial n_{e}}{\partial\mu_{e}}\right)_{T}+\left(\frac{\partial n_{h}}{\partial\mu_{h}}\right)_{T}\right]^{-1}.$ (S10)
One can see that at the point $\mu=0$ the number of electrons coincides with the number of holes ($n_{e}=n_{h}$), i.e. the first multiplier in Eq. (S10) turns zero, while with use of Eq. (S6) it is easy to find that the second multiplier is equal to $\pi v_{F}/2$. Hence, the entropy per particle and Seebeck coefficient in the Dirac point turn out to be equal to zero:
@@
III S3: Derivation of the expression for conductivity
Here we provide the detailed calculation of the integral for conductivity:
$\sigma=\frac{{{e^{2}}}}{{8hT}}\int\limits_{-\infty}^{+\infty}{\left[{3+\frac{2}{\pi}\arctan\left({\left({E-{E_{p}}}\right)\tau}\right)}\right]\frac{{dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}.$ (S11)
We split integral Eq. (S11) into parts $\sigma={\sigma_{1}}+{\sigma_{2}}$, where
${\sigma_{1}}=\frac{{3{e^{2}}}}{{8hT}}\int\limits_{-\infty}^{+\infty}{\frac{{dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}=\frac{{3{e^{2}}}}{{2h}},$ (S12)
@@ III S3: Derivation of the expression for conductivity
${\sigma_{2}}=\frac{{{e^{2}}}}{{4\pi hT}}\int\limits_{-\infty}^{+\infty}{\frac{{\arctan\left({\left({E-{E_{p}}}\right)\tau}\right)dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}.$ (S13)
Applying integration by parts to Eq. (S13) one can rewrite the integral as
${\sigma_{2}}=-\frac{{{e^{2}}T\tau}}{{\pi h}}\int\limits_{-\infty}^{+\infty}{\frac{{\tanh zdz}}{{1+{{\left({2T\tau}\right)}^{2}}{{\left({z-\frac{\left(E_{p}-\mu\right)}{2T}}\right)}^{2}}}}},$ (S14)
The latter integral can be evaluated using the Cauchy’s theorem by means of residues. The appearing summation over the poles of $\tanh z$ can be performed in terms of digamma-function and after straightforward calculations one finds:
${\sigma_{2}}=-\frac{{{e^{2}}}}{{\pi h}}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}+\frac{1}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right),$ (S15)
where the symmetry property of the digamma function $\psi\left({\bar{z}}\right)=\overline{\psi\left(z\right)}$ was applied [1].
Therefore, the final result for conductivity is expressed as
$\sigma\!=\sigma^{(sm)}\!=\!\frac{{{e^{2}}}}{{2h}}\left[{3\!-\!\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}\!+\!\frac{1}{{2\pi T\tau}}\!+\!i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}\right].$ (S16)
In the case of $\tau\to\infty$ using the relation between the imaginary part of the digamma function and the hyperbolic tangent [1] one can reduce Eq. (S15) to the expression
${\sigma_{2}}\!=\!-\frac{{{e^{2}}}}{{\pi h}}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}\!+\!i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)=\!-\frac{{{e^{2}}}}{{2h}}\tanh\left({\frac{{{E_{p}}\!-\!\mu}}{{2T}}}\right).$ (S17)
and reproduce Eq. (15) in the main text of the paper.

### v2

@@ IV.2 A parabolic subband: Mott’s formula
${\sigma^{(sm)}}\!=\!\frac{{{e^{2}}}}{{4h}}\frac{1}{{4T}}\int_{-\infty}^{\infty}{\left[{3\!+\!\frac{2}{\pi}\arctan\left({\frac{{\left({E\!-\!{E_{p}}}\right)\tau}}{\hbar}}\right)}\right]}\frac{{dE}}{{{{\cosh}^{2}}\frac{{E\!-\!\mu}}{{2T}}}}.$ (23)
Integration of the second term by parts followed by the application of the Cauchy’s theorem (see Appendix C) yields the conductance explicitly:
$\sigma^{(sm)}\!=\!\frac{{{e^{2}}}}{{4h}}\left[{3\!-\!\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}\!+\!\frac{\hbar}{{2\pi T\tau}}\!+\!i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}\right].$ (24)
@@
$S_{M}^{\left({{{sm}}}\right)}(\mu)=\frac{1}{{3e}}\,\frac{{{\mathop{\rm Re}\nolimits}{\psi^{\prime}}\left({\frac{1}{2}+\frac{\hbar}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}}{{3-\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}+\frac{\hbar}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}}.$ (25)
@@
In the limiting cases where the chemical potential is either close to the bottom of the parabolic subband or sufficiently far from it one finds
$S_{M}^{(sm)}=\frac{1}{{3e}}\begin{cases}\frac{1}{3}{\mkern 1.0mu}{\psi^{\prime}}\left({\frac{1}{2}\!+\!\frac{\hbar}{{2\pi T\tau}}}\right)\!+\!\left[\psi^{\prime}\left({\frac{1}{2}+\frac{\hbar}{{2\pi T\tau}}}\right)\right]^{2}\cdot\frac{{{E_{p}}-\mu}}{72\pi^{2}T}\\ \qquad\qquad\qquad\qquad\text{when}\qquad|E_{p}-\mu|\ll T,\\ \frac{{32\pi\hbar}}{11\cdot T\tau}\frac{{T^{2}}}{{{{\left({{E_{p}}-\mu}\right)}^{2}}}}\qquad\text{when}\qquad|E_{p}-\mu|\gg T.\end{cases}.$
@@
Figure 4 (a,b) shows the results of calculation of the Seebeck coefficient as a function of the chemical potential in the vicinity of the bottom of the second electronic subband (j=2) performed accounting for the smearing of DOS. The curves in the panels (a) and (b) are calculated with use of the Kelvin’s and Mott’s approaches, respectively. One can see that even in the case of a strong smearing Kelvin’s and Mott’s approaches yield qualitatively different results, especially in what concerns the change of sign of the Seebeck coefficient in the vicinity of the topological phase transition point.
V Conclusions
@@
The ballistic regime offers peculiar modifications of some well-known thermoelectric relations. In particular, it turns out that Mott’s formula for the Seebeck effect is incorrect for one-dimensional parabolic bands, while the Kelvin’s formula remains fully accurate in this case. This is important also because the application of Kelvin’s formula is less challenging numerically than the application of the Mott’s formula in many cases. The differences between two approaches become apparent in the vicinity of resonances between the electronic chemical potential and the bottoms of energy subbands characterised by parabolic dispersion. Mott’s formula predicts a finite value of the Seebeck coefficient at the resonance and no change of sign. In contrast, the Kelvin’s formula predicts divergence and the change of sign of the Seebeck coefficient at the resonance between the chemical potential and the bottom of a parabolic subband. The latter result is characteristic of most of topological phase transitions that occur once a new energy subband comes into play. Our analysis shows that both Kelvin’s and Mott’s expressions predict zero contributions to the Seebeck coefficient from the linear dispersion band in metallic CNTs in the ballistic regime. In contrast, the contribution of parabolic bands is not zero even far from the resonance points, according to the Kelvin’s (but not Mott’s!) formula. This conclusion can be easily verified experimentally. Any deviation of the Seebeck coefficient of CNTs from zero in the ballistic regime would characterize the inaccuracy of Mott’s formula. Finally, we note that in 2D and 3D cases, the Kelvin’s formula also fails in the ballistic regime. This is because counter-propagating non-dissipative currents of ”hot” and ”cold” electrons are formed due to the combined actions of the temperature and voltage drops in these cases. The total electric current remains zero, but the system cannot be described by a single electro-chemical potential. We are looking forward for the experimental manifestations of these theoretical results.
@@
Appendix A The link between the temperature derivative of the chemical potential and the density of states
We consider the electronic contribution to the Seebeck effect. More general relations accounting for the hole contribution too are given in the Appendix B. To start with, with no loss of generality, the temperature derivative of the chemical potential can be expressed as follows:
$\left(\frac{\partial\mu}{\partial T}\right)_{n}=\left(\frac{\partial n}{\partial T}\right)_{\mu}\left(\frac{\partial n}{\partial\mu}\right)_{T}^{-1}.$ (26)
@@ Appendix A The link between the temperature derivative of the chemical potential and the density of states
$n\left(\mu,T\right)=\int\limits_{0}^{+\infty}\frac{\nu(E)dE}{\exp\left(\frac{E-\mu}{T}\right)+1}.$ (27)
Required derivatives can be performed assuming $T\ll\mu$
$\left(\frac{\partial n}{\partial T}\right)_{\mu}=\frac{\pi^{2}T}{3}\frac{d\nu\left(\mu\right)}{d\mu}+\frac{7\pi^{4}T^{3}}{90}\frac{d^{3}\nu(\mu)}{d\mu^{3}}+\mathcal{O}\left(\frac{T^{5}}{\mu^{5}}\right),$ (28)
@@
$\left(\frac{\partial n}{\partial\mu}\right)_{T}=\nu\left(\mu\right)+\frac{\pi^{2}T^{2}}{6}\frac{d^{2}\nu(\mu)}{d\mu^{2}}+\mathcal{O}\left(\frac{T^{4}}{\mu^{4}}\right).$ (29)
Hence, in the low temperature limit
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)_{N,V}=\frac{\pi^{2}}{3e}T\frac{d\ln\nu(\mu)}{d\mu}.$ (30)
Appendix B The vicinity of the Dirac point
@@
$n_{e,h}\left(\mu,T\right)=\frac{2}{\pi\hbar v_{F}}\int\limits_{0}^{+\infty}\frac{dE}{\exp\left(\frac{E-\mu_{e,h}}{T}\right)+1},$ (31)
@@
$n_{e}-n_{h}=n_{0}.$ (32)
Here $-|e|n_{0}$ is the overall charge density. Substituting Eq. (31) in Eq. (32) and performing integration one finds
$\mu=\mu_{e}=\frac{\pi\hbar v_{F}n_{0}}{2}.$ (33)
@@
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)_{n}=\frac{1}{e}\left(\frac{\partial n_{0}}{\partial T}\right)_{\mu}\left(\frac{\partial n_{0}}{\partial\mu}\right)_{T}^{-1}=0.$ (34)
B.1 The special point $\mu=0$
@@
$\left[\left(\frac{\partial n_{e}}{\partial T}\right)_{\mu_{e}}-\left(\frac{\partial n_{h}}{\partial T}\right)_{\mu_{h}}\right]\left[\left(\frac{\partial n_{e}}{\partial\mu_{e}}\right)_{T}+\left(\frac{\partial n_{h}}{\partial\mu_{h}}\right)_{T}\right]^{-1}.$ (35)
One can see that at the point $\mu=0$ the number of electrons coincides with the number of holes ($n_{e}=n_{h}$), i.e. the first multiplier in Eq. (35) turns zero, while with use of Eq. (31) it is easy to find that the second multiplier is equal to $\pi v_{F}/2$. Hence, the entropy per particle and Seebeck coefficient in the Dirac point turn out to be equal to zero:
@@
Appendix C Derivation of the expression for conductance
Here we provide the detailed calculation of the integral for conductance:
$\sigma=\frac{{{e^{2}}}}{{8hT}}\int\limits_{-\infty}^{+\infty}{\left[{3+\frac{2}{\pi}\arctan\left({\left({E-{E_{p}}}\right)\hbar^{-1}\tau}\right)}\right]\frac{{dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}.$ (36)
We split integral Eq. (36) into parts $\sigma={\sigma_{1}}+{\sigma_{2}}$, where
${\sigma_{1}}=\frac{{3{e^{2}}}}{{8hT}}\int\limits_{-\infty}^{+\infty}{\frac{{dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}=\frac{{3{e^{2}}}}{{2h}},$ (37)
@@ Appendix C Derivation of the expression for conductance
${\sigma_{2}}=\frac{{{e^{2}}}}{{4\pi hT}}\int\limits_{-\infty}^{+\infty}{\frac{{\arctan\left({\left({E-{E_{p}}}\right)\hbar^{-1}\tau}\right)dE}}{{{{\cosh}^{2}}\frac{{E-\mu}}{{2T}}}}}.$ (38)
Applying integration by parts to Eq. (38) one can rewrite the integral as
${\sigma_{2}}=-\frac{{{e^{2}}T\tau}}{{\pi h}}\int\limits_{-\infty}^{+\infty}{\frac{{\tanh zdz}}{{1+{{\left({2T\hbar^{-1}\tau}\right)}^{2}}{{\left({z-\frac{\left(E_{p}-\mu\right)}{2T}}\right)}^{2}}}}},$ (39)
where we introduced new variable $z=\frac{{\left({E-{E_{p}}}\right)\hbar^{-1}\tau+\delta}}{{2T\tau}}$ with the parameter $\delta=\left({{E_{p}}-\mu}\right)\hbar^{-1}\tau$. The latter integral can be evaluated using the Cauchy’s theorem by means of residues. The appearing summation over the poles of $\tanh z$ can be performed in terms of digamma-function and after straightforward calculations one finds:
${\sigma_{2}}=-\frac{{{e^{2}}}}{{\pi h}}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}+\frac{\hbar}{{2\pi T\tau}}+i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right),$ (40)
where the symmetry property of the digamma function $\psi\left({\bar{z}}\right)=\overline{\psi\left(z\right)}$ was applied [32]. The similar structure of the expression for conductance in terms of the polygamma function was also obtained in Refs. 34, 35, 36, 37.
Therefore, the final result for conducance is expressed as
$\sigma\!=\sigma^{(sm)}\!=\!\frac{{{e^{2}}}}{{2h}}\left[{3\!-\!\frac{2}{\pi}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}\!+\!\frac{\hbar}{{2\pi T\tau}}\!+\!i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)}\right].$ (41)
In the case of $\tau\to\infty$ using the relation between the imaginary part of the digamma function and the hyperbolic tangent [32] one can reduce Eq. (40) to the expression
${\sigma_{2}}\!=\!-\frac{{{e^{2}}}}{{\pi h}}{\mathop{\rm Im}\nolimits}\psi\left({\frac{1}{2}\!+\!i\frac{{{E_{p}}-\mu}}{{2\pi T}}}\right)=\!-\frac{{{e^{2}}}}{{2h}}\tanh\left({\frac{{{E_{p}}\!-\!\mu}}{{2T}}}\right).$ (42)
and reproduce Eq. (13).